PS1 IPP Czar Logs for the week 2010.12.27 - 2011.01.02

(Up to PS1 IPP Czar Logs)

Monday : 2010.12.27

11:55 (heather) Processing is done, except for a few faults. Gene debugged and pushed the nights data through burntool (at 4:30am and at 8am). Neither of the holiday czars are sure how to fix burntool if there are more problems. Heather notes that isp is not being registered, and this is because it is commented out in registration, pending burntool testing and fixing.

11:57 (heather) Heather is running a stdscience like pantasks from the current trunk on c14, to do tests of the new static masks.

17:00 (heather) Heather added MD01.jtrp (reprocessing MD fields that JT found that were not completed before), also, Gene added STS.20101202.

Tuesday : 2010.12.28

6:59 (roy) restarted summitcopy, which crashed at 23:27:35 last night.

8:30 (heather) Heather noticed that registration was not complete. registration's revert is failing:

regtool -revertprocessedimfile -dbname gpc1 -limit 128

because it doesn't understand -limit. Heather reverted by hand:

regtool -revertprocessedimfile -dbname gpc1 -exp_id_begin 128000

and the unregistered files registered.

17:00 (heather) heather's pantasks on c14 is running magictest.% labels (for testing the static mask). The additional labels in the regular pantasks (Gene's STS and heather's MD01) are still processing.

18:04 (heather) roboczar said stdscience was down, heather investigated, it's been running for a while!?

Wednesday : 2010.12.29

7:50 (roy) No data last night, remaining exposures from yesterday were completed by midnight.

11:01 (heather) magictest labels still warping. MD01.jtrp now stacking

16:26 (heather) magictest labels - 3pi and oss are diffing, rest still warping. MD02.jtrp and MD03.jtrp added into regular stdscience

Thursday : 2010.12.30

8:00 (roy) last night's exposures are not registered, but summitcopy is running.

9:20 (heather) heather did the following:

regtool -revertprocessedimfile -dbname gpc1 -exp_id_begin 128000

This reverted 61 imfiles.

and, before heather noticed that registration was working, restarted the registration pantasks. It's working now, but has a long ways to go before we can start processing data today.

14:30 (heather and roy) noticed burntool/registration is stuck. reverts, restarting pantasks didn't work. running by hand worked.

16:46 (heather) only looks for things to burntool for the current day. Since the day cycled over, there was nothing more to do. Heather hacked '2010-12-30' into $today in, recompiled, restarted pantasks, and things are registering now.

17:27 (heather) reverted to the regular one (the one that grabs the burns for the current day), restarted registration pantasks just to make sure there was nothing funny, and typed 'run' in heather's stdsci and the ipp stdscience.

Friday : 2010.12.31

RCUH holiday.

Saturday : 2011.01.01

Heather reverted (using regtool revert) burntool/registration.

Sunday : 2011.01.02

heather reverted (using regtool revert) burntool/registration